OpenStudy (anonymous):

Write a polynomial function f of least degree that has rational coefficients, a leading coefficient of 1, and the given zeros. 1) -2, -1, 1

OpenStudy (anonymous):

start with the zeros... y = (x+2)(x+1)(x-1) multiply out that right side....

OpenStudy (anonymous):

Reverse solve for x for the zeroes. x = -2 x + 2 = 0 x = -1 x + 1 = 0 x = 1 x - 1 = 0 f(x) = (x + 2)(x + 1)(x - 1) Can you multiply it out?

OpenStudy (anonymous):

f(x) = (x+2)(x+1)(x-1) Multiply it out See that putting -1,-1 or 1 in the abuv function gives 0

OpenStudy (anonymous):

what about when im given i, 2 - √3